# Data-based control of Logical Networks

### v1

@@ 4 Data-driven safe control
Conclusion: if $\mathcal{S}^{*}=\mathcal{L}_{N}$, then return (‘Yes’, $\mathcal{S}^{*}$)
                else, return (‘No’, $\mathcal{S}^{*}$)
@@
Theorem 13.
Problem 2 is solvable (i.e., the data $({\bf x}_{d},{\bf u}_{d})$ are informative for safe control with respect to the unsafe set ${\mathcal{X}}_{u}$) if and only if the following two conditions hold:
  • (i)
    for every $\delta_{N}^{j}\in{\mathcal{X}}_{s}$ there exist $k\in[1,T]$ and $\delta_{N}^{\ell}\in{\mathcal{X}}_{s}$ such that (15) holds;
  • (ii)
    ${\mathcal{S}^{*}}=\bigcup_{{\bf x}\in\mathcal{X}_{s}}\mathcal{S}({\bf x})=\mathcal{L}_{N}$.
Moreover, the generic $r$-th column of a possible feedback matrix $K\in{\mathcal{L}}_{M\times N}$ is given by  ${\rm col}_{r}(K)={\bf u}_{r},\forall r\in[1,N],$ where ${\bf u}_{r}$ is determined through Algorithm 3.
Proof.
(Necessity). Assume, by contradiction, that condition (i) does not hold. This implies that there is at least one state in $\mathcal{X}_{s}$ whose immediate successors, as they can be deduced from data, are all in $\mathcal{X}_{u}$. But this means that there exists at least one BCN in $\mathcal{B}_{d}$ for which condition (i) in Proposition 12 does not hold. For that BCN the safe control problem is not solvable.
Assume now that (i) holds, but (ii) does not, namely $\mathcal{S}^{*}\neq\mathcal{L}_{N}$. We want to construct a BCN in ${\mathcal{B}}_{d}$ for which one of the states in $\mathcal{X}_{u}$ cannot reach a safe state in a finite number of steps. This requires to impose all the state transitions that are not uniquely identified by the data. If there exists $\delta^{r}_{N}\in{\mathcal{X}}_{u}$ that does not appear as a column of $X_{p}$, then for every $i\in[1,M]$ we impose $\tilde{L}_{i}\delta^{r}_{N}=\delta^{r}_{N}$. On the other hand, if $\delta^{r}_{N}=X_{p}\delta^{k}_{T}$ for some $k\in[1,T]$, we define $\mathcal{I}_{r}$ the set of indices $i\in[1,M]$ such that $U_{p}\delta^{k}_{T}=\delta^{i}_{M}$. If $\mathcal{I}_{r}\subsetneq[1,M]$, we impose $\tilde{L}_{i}\delta^{r}_{N}=\delta^{r}_{N}$ for every $i\not\in\mathcal{I}_{r}$. Clearly, $\mathcal{X}_{s}$ is not reachable from $\delta^{r}_{N}\in{\mathcal{X}}_{u}$ and hence by Proposition 12, the safe control problem for such BCN in ${\mathcal{B}}_{d}$ is not solvable.
(Sufficiency). It is a direct consequence of Proposition 12 and subsequent comments.
The second part follows trivially from Algorithm 3. ∎
Example 14, below, explores on a small size network the results of this section and of the previous one pertaining data-driven state-feedback stabilization and safe control.
@@
$L$ $=$ $\left[\begin{array}[]{c|c|c}L_{1}&L_{2}&L_{3}\end{array}\right]=\big{[}\delta_{7}^{4}\ \delta_{7}^{2}\ \delta_{7}^{2}\ \delta_{7}^{5}\ \delta_{7}^{2}\ \delta_{7}^{7}\ \delta_{7}^{5}\ |$
@@
The equilibrium points of this BCN are $\{\delta_{7}^{2},\delta_{7}^{3},\delta_{7}^{6},\delta_{7}^{7}\}$. We perform an offline experiment in the time interval $[0,T]$, with $T=12$ and we collect the following data:
@@
$\tilde{L}$ $=$ $\left[\begin{array}[]{c|c|c}\tilde{L}_{1}&\tilde{L}_{2}&\tilde{L}_{3}\end{array}\right]=\big{[}\delta_{7}^{4}\ \delta_{7}^{2}\ \delta_{7}^{2}\ *\ *\ *\ *\ |$
@@
where the symbol * stands for an arbitrary vector (in ${\mathcal{L}}_{7}$). By applying Algorithm 1, we obtain that the set of the equilibrium points compatible with all the BCNs in $\mathcal{B}_{d}$ is $\mathcal{X}_{e}^{d}=\{\delta_{7}^{2},\delta_{7}^{3},\delta_{7}^{7}\}.$ We now check if there exists at least one state in $\mathcal{X}_{e}^{d}$ that is globally reachable. To this end, we follow the procedure outlined in Algorithm 2.
$\bullet$  For ${\bf x}_{e}^{1}\triangleq\delta_{7}^{2}$, we get:
$\mathcal{S}_{0}=\{\delta_{7}^{2}\}$, $\mathcal{S}_{1}=\{\delta_{7}^{3}\}$, $\mathcal{S}_{2}=\{\delta_{7}^{4}\}$, $\mathcal{S}_{3}=\{\delta_{7}^{1}\}$, $\mathcal{S}_{4}=\{\delta_{7}^{5}\}$, $\mathcal{S}_{5}=\{\delta_{7}^{6}\}$, $\mathcal{S}_{6}=\{\delta_{7}^{7}\}$, and hence $\mathcal{S}({\bf x}_{e}^{1})={\mathcal{L}}_{7}$.
$\bullet$   For ${\bf x}_{e}^{2}\triangleq\delta_{7}^{3}$, we get:
$\mathcal{S}_{0}=\{\delta_{7}^{3}\}$, $\mathcal{S}_{1}=\{\delta_{7}^{4}\}$, $\mathcal{S}_{2}=\{\delta_{7}^{1}\}$, $\mathcal{S}_{3}=\{\delta_{7}^{5}\}$, $\mathcal{S}_{4}=\{\delta_{7}^{6}\}$, $\mathcal{S}_{5}=\{\delta_{7}^{7}\}$, $\mathcal{S}_{6}=\emptyset$, and hence $\mathcal{S}({\bf x}_{e}^{2})\neq{\mathcal{L}}_{7}$.
$\bullet$   For ${\bf x}_{e}^{2}\triangleq\delta_{7}^{7}$, we get:
$\mathcal{S}_{0}=\{\delta_{7}^{7}\}$, $\mathcal{S}_{1}=\{\delta_{7}^{1}\}$, $\mathcal{S}_{2}=\{\delta_{7}^{5}\}$, $\mathcal{S}_{3}=\{\delta_{7}^{6}\}$, $\mathcal{S}_{4}=\emptyset$, and hence $\mathcal{S}({\bf x}_{e}^{2})\neq{\mathcal{L}}_{5}$.
Therefore, by Theorem 10, the data are informative for stabilization to ${\bf x}_{e}^{1}$ by state feedback. The state feedback matrix $K=\begin{bmatrix}\delta_{3}^{1}&\delta_{3}^{1}&\delta_{3}^{1}&\delta_{3}^{3}&\delta_{3}^{3}&\delta_{3}^{2}&\delta_{3}^{3}\end{bmatrix}\in\mathcal{L}_{3\times 7}$ is obtained through Algorithm 2.
Assume now that we want to implement a state feedback safe control with respect to the unsafe set $\mathcal{X}_{u}=\{\delta_{7}^{3},\delta_{7}^{4},\delta_{7}^{7}\}$. We need to understand if the data are informative for safe control with respect to $\mathcal{X}_{u}$, or, equivalently, if conditions (i) and (ii) of Theorem 13 are satisfied. In this case $\mathcal{X}_{s}=\{\delta_{7}^{1},\delta_{7}^{2},\delta_{7}^{5},\delta_{7}^{6}\}$, and hence it is easy to check on data that for every state in $\mathcal{X}_{s}$ there is a transition ending in a state that is still inside $\mathcal{X}_{s}$, namely condition (i) of Theorem 13 holds. Now we apply Algorithm 3 to verify whether $\mathcal{X}_{s}$ is globally reachable in every BCN in ${\mathcal{B}}_{d}$. We obtain: $\mathcal{S}_{0}=\mathcal{X}_{s}$, $\mathcal{S}_{1}=\{\delta_{7}^{3},\delta_{7}^{7}\}$, and $\mathcal{S}_{2}=\{\delta_{7}^{4}\}$, and hence $\mathcal{S}^{*}=\mathcal{L}_{7}$. This means that also condition (ii) of Theorem 13 is satisfied, and thus the data are informative for safe control with respect to $\mathcal{X}_{u}$. Algorithm 3 provides the state feedback matrix $K=\begin{bmatrix}\delta_{3}^{2}&\delta_{3}^{1}&\delta_{3}^{1}&\delta_{3}^{3}&\delta_{3}^{3}&\delta_{3}^{2}&\delta_{3}^{3}\end{bmatrix}$.
@@ 5 Data-driven output regulation
As before, we assume to collect some input and state data on a finite time interval $[0,T],T\in\mathbb{Z}_{+}$. However, this time we collect also output trajectories. Let ${\bf y}_{d}\triangleq\{{\bf y}_{d}(t)\}_{t=0}^{T-1}$ be the corresponding output sequence. We group the output data into a matrix analogous to the ones in (9)$\div$(11):
$Y_{p}\triangleq\begin{bmatrix}{\bf y}_{d}(0)&{\bf y}_{d}(1)&\dots&{\bf y}_{d}(T-1)\end{bmatrix}\in\mathcal{L}_{P\times T}.$ (22)
We extend the concept of compatibility of a BCN with the collected data to the case in which the BCN is described by (4), i.e., we include the output equation, and the collected data are $({\bf x}_{d},{\bf u}_{d},{\bf y}_{d})$. The set of Boolean control networks (4) compatible with $({\bf x}_{d},{\bf u}_{d},{\bf y}_{d})$ is
$\!\!\!\!\mathcal{B}_{d}^{o}\!$$\triangleq\{(\tilde{L},\tilde{H}),\tilde{L}\in\mathcal{L}_{N\times NM},\tilde{H}\in\mathcal{L}_{P\times N}:\forall\ t\in[0,T-1]$
$\!\!\!{\bf x}_{d}(t+1)=\tilde{L}\ltimes{\bf u}_{d}(t)\ltimes{\bf x}_{d}(t),\ {\bf y}_{d}(t)=\tilde{H}{\bf x}_{d}(t)\}.\!\!\!$ (23)
Obviously, the pair $(L,H)$ corresponding to the BCN (4) that generated the data belongs to $\mathcal{B}_{d}^{o}$. Accordingly, the notion of informativity for identifiability has to be extended.
@@
The data $({\bf x}_{d},{\bf u}_{d},{\bf y}_{d})$ are called informative for identifiability of the BCN (4) if ${\mathcal{B}}_{d}^{o}\equiv\{(L,H)\}$.
Even if the definition is different, the characterization of the new notion of informativity for identifiability is still the one in Lemma 6: the matrix in (13) must be of full row rank. Necessity is obvious. On the other hand, if the matrix in (13) is of full row rank, then we deduce from the collected data every possible state transition that can occur in the network. But this implies that all possible states appear at least once in $X_{p}$ (and consequently also the corresponding output in $Y_{p}$). This allows to uniquely identify also the matrix $H$ in (4b), and hence the data $({\bf x}_{d},{\bf u}_{d},{\bf y}_{d})$ are informative for identifiability of the whole BCN (4).
In the sequel, we assume again that the data are not informative for the identifiability of the BCN.
Problem 3.
Given the data $({\bf x}_{d},{\bf u}_{d},{\bf y}_{d})$, and some desired value ${\bf y}^{*}\in{\mathcal{L}}_{P}$ for the output in (4b), determine (if possible) a state feedback matrix $K\in\mathcal{L}_{M\times N}$ such that, by applying ${\bf u}(t)=K{\bf x}(t)$, for every initial condition ${\bf x}(0)$, the output trajectories of all the BCNs compatible with the data satisfy ${\bf y}(t)={\bf y}^{*}$ for every $t\geq\tau,\exists\tau\in\mathbb{Z}_{+}$.
When Problem 3 is solvable, we will say that the data $({\bf x}_{d},{\bf u}_{d},{\bf y}_{d})$ are informative for output regulation to ${\bf y}^{*}\in{\mathcal{L}}_{P}$ by state feedback. In order to solve this problem, we introduce some definitions and results, taken from the model-based literature on BNs and BCNs (see, e.g., Fornasini & Valcher (2013a, 2014)).
Definition 16.
@@
Given a BCN (4), the output regulation problem to ${\bf y}^{*}\in{\mathcal{L}}_{P}$ by state feedback is said to be solvable if there exist $K\in\mathcal{L}_{M\times N}$ and $\tau\in\mathbb{Z}_{+}$ such that for every ${\bf x}(0)\in\mathcal{L}_{N}$, ${\bf y}(t)={\bf y}^{*},\forall t\geq\tau$, when ${\bf u}(t)=K{\bf x}(t).$
@@
The first step of the model-based solution to this problem consists in finding all the states that generate the desired output ${\bf y}^{*}$, namely the states belonging to the set $\mathcal{X}({\bf y}^{*})\triangleq\{\delta_{N}^{j}:H\delta_{N}^{j}={\bf y}^{*}\}$. Once we have determined this set, we have to ensure that, under some state feedback law, the state evolution of the resulting BN (8) is constrained within $\mathcal{X}({\bf y}^{*})$, from some time $\tau\in\mathbb{Z}_{+}$ onwards, which imposes that $\mathcal{X}({\bf y}^{*})$ contains either an equilibrium point of the BN or a limit cycle (whose definition is given below).
Definition 17.
Fornasini & Valcher (2013b) Given a BN as in (3), an ordered sequence of distinct logical vectors $(\delta_{N}^{i_{1}},\delta_{N}^{i_{2}},\dots,\delta_{N}^{i_{k}})$ is a limit cycle $\mathcal{C}$ if, taken ${\bf x}(0)=\delta_{N}^{i_{l}}$ for some $l\in[1,k]$, the corresponding state evolution ${\bf x}(t)$ is periodic of period $k$ and, for every $t\in\mathbb{Z}_{+}$, ${\bf x}(t)=\delta_{N}^{i_{j}}$, with $j\in[1,k]$ satisfying $j=(t+l){\rm mod}k$.
@@
Note that an equilibrium point is a limit cycle of unitary length. We denote by $\mathcal{C}({\bf y}^{*})$ the set of all limit cycles contained in $\mathcal{X}({\bf y}^{*})$. Once we have ascertained that $\mathcal{C}({\bf y}^{*})\neq\emptyset$, the final step consists in verifying that it is globally reachable, meaning that the union of the basins of attraction of the states in $\mathcal{C}({\bf y}^{*})$ covers the whole state space $\mathcal{L}_{N}$.
@@
In the following proposition we state necessary and sufficient conditions for solving the output regulation problem by state feedback in a model-based set-up.
@@
We can now provide the solution to Problem 3. The first step consists in finding, from data, the set of states that generate the desired output ${\bf y}^{*}$ in all BCNs in ${\mathcal{B}}_{d}^{o}$, i.e.:
$\mathcal{X}^{d}({\bf y}^{*})\!=\!\left\{\delta_{N}^{j}\in{\mathcal{L}}_{N}:\exists k\in[1,T]\ \text{s.t.}\begin{bmatrix}X_{p}\cr Y_{p}\end{bmatrix}\delta_{N}^{k}=\!\begin{bmatrix}\delta_{N}^{j}\cr{\bf y}^{*}\end{bmatrix}\right\}\!.$
To identify this set, we can proceed as follows.
Algorithm 4: Construction of $\mathcal{X}^{d}({\bf y}^{*})$
@@
Output: The set $\mathcal{X}^{d}({\bf y}^{*})$ of all states of the BCNs in ${\mathcal{B}}_{d}^{o}$ that generate the desired output ${\bf y}^{*}$.
Initialization:
Set $\mathcal{X}^{d}({\bf y}^{*})=\emptyset$.
Procedure:

### v2

@@ 4 Reachability property
Conclusion: Return ($({\bf u}_{1},{\bf u}_{2},\dots,{\bf u}_{N})$, $\mathcal{S}^{*}$)
Note that Algorithm 1 identifies first the set ${\mathcal{S}}_{1}$ of states in ${\mathcal{L}}_{N}\setminus\mathcal{X}$ that have a successor in $\mathcal{S}_{0}=\mathcal{X}$. For each such state, say $\delta^{r}_{N}$, it memorizes in ${\bf u}_{r}$ one of the values of the input that allows the transition from $\delta^{r}_{N}$ to one of the states in $\mathcal{X}$, say $\delta^{q}_{N}$ (referring to the algorithm notation). Then it identifies the set ${\mathcal{S}}_{2}$ of the states that do not belong to ${\mathcal{S}}_{0}\cup{\mathcal{S}}_{1}=\color[rgb]{0,0,0}{\mathcal{X}}\cup{\mathcal{S}}_{1}$, but have a successor in ${\mathcal{S}}_{1}$. Again, for each $\delta^{r}_{N}\in{\mathcal{S}}_{2}$, it memorizes one of the values of the input that allows the transition from $\delta^{r}_{N}$ to some state in ${\mathcal{S}}_{1}$. Since the set of states of a BCN is finite, there exists $d\geq 1$ such that ${\mathcal{S}}_{d}=\emptyset$. If at that stage $\bigcup_{i=0}^{d-1}{\mathcal{S}}_{i}=\mathcal{L}_{N}$, this means that $\mathcal{X}$ is globally reachable.
5 Equilibrium points of a BCN
We now recall the notions of equilibrium point and limit cycle in the context of logical systems.
Definition 8.
Note that ${\bf x}_{e}=\delta^{j}_{N}$ is an equilibrium point of a BN or a BCN if and only if the node $j$ has a self-loop in ${\mathcal{D}}(L)$ or ${\mathcal{D}}(L_{tot})$, respectively. This amounts to saying that $L]_{jj}$ or $L_{tot}]_{jj}$, respectively, is unitary.
Based on the comments about $L_{tot}$ in the previous section, it is immediate to deduce that the set of equilibrium points compatible with the available data, say, $\mathcal{X}_{e}^{d}$, namely the set of states that are equilibria in all the BCNs in $\mathcal{B}_{d}$, coincides with the set of states $\delta^{j}_{N}$ such that $L_{tot}^{d}]_{jj}=1$, or, equivalently, $X_{f}(X_{p})^{\top}]_{jj}>0$, which is consistent with the analysis proposed in Section 3.2 of Li et al. (2025). As an alternative, we can notice that $L_{tot}^{d}]_{jj}=1$ if and only if $\exists k\in[1,T]$ such that
$\begin{bmatrix}X_{p}\cr X_{f}\end{bmatrix}\delta_{T}^{k}=\begin{bmatrix}\delta_{N}^{j}\cr\delta_{N}^{j}\end{bmatrix}\Leftrightarrow(X_{p}\odot X_{f})\delta^{k}_{T}=\delta_{N}^{j},$
where we recall that $X_{p}\odot X_{f}$ is the Hadamard product of $X_{p}$ and $X_{f}$. This characterization has the advantage of allowing us to identify the input to which the equilibrium point corresponds as ${\bf u}_{j}\triangleq U_{p}\delta_{T}^{k}=\delta_{M}^{i},\exists i\in[1,M]$.
Since the number of possible states is finite, as time evolves, the state of a Boolean network either becomes constant (and hence reaches an equilibrium point) or it becomes periodic. In the latter case, we talk about a limit cycle. Note that an equilibrium point is actually a limit cycle of unit length.
Definition 9.
Fornasini & Valcher (2013b) Given a BN as in (5), an ordered sequence of distinct logical vectors $(\delta_{N}^{i_{1}},\delta_{N}^{i_{2}},\dots,\delta_{N}^{i_{k}})$ is a limit cycle $\mathcal{C}$ if, taken ${\bf x}(0)=\delta_{N}^{i_{\ell}}$ for some $\ell\in[1,k]$, the corresponding state evolution ${\bf x}(t)$ is periodic of period $k$ and, for every $t\in\mathbb{Z}_{+}$, ${\bf x}(t)=\delta_{N}^{i_{j}}$, with $j\in[1,k]$ satisfying $j=(t+\ell){\rm mod}k$.
We will make use of this concept in Section 7.
6 Data-driven safe control
In Fornasini & Valcher (2016) a complete characterization of the problem solvability has been provided.
Proposition 10.
  • i)
  • ii)
Remark 11.
It is worth remarking that safe control to a set ${\mathcal{X}}_{s}$ is similar to the problem of stabilizing a BCN to the set ${\mathcal{X}}_{s}$, but it does not coincide with it. Indeed, while stabilization only requires that every state trajectory eventually enters ${\mathcal{X}}_{s}$, safe control imposes the additional constraint that states belonging to ${\mathcal{X}}_{s}$ can never leave ${\mathcal{X}}_{s}$, not even in a transient phase.
Problem 1.
When Problem 1 is solvable, the data $({\bf x}_{d}$, ${\bf u}_{d})$ will be called informative for safe control (with respect to the unsafe set ${\mathcal{X}}_{u}$). In the following theorem we provide necessary and sufficient conditions for the solvability of Problem 1.
Theorem 12.
Consider the unsafe set ${\mathcal{X}}_{u}$ and set ${\mathcal{X}}_{s}={\mathcal{L}}_{N}\setminus{\mathcal{X}}_{u}=\{\delta^{i_{1}}_{N},\delta^{i_{2}}_{N},\dots,\delta^{i_{d}}_{N}\}$. Problem 1 is solvable (i.e., the data $({\bf x}_{d},{\bf u}_{d})$ are informative for safe control with respect to the unsafe set ${\mathcal{X}}_{u}$) if and only if the following two conditions hold:
  • i)
    for every $\delta_{N}^{j}\in{\mathcal{X}}_{s}$ there exist $k\in[1,T]$ and $\delta_{N}^{\ell}\in{\mathcal{X}}_{s}$ such that
    $\begin{bmatrix}X_{p}\cr X_{f}\end{bmatrix}\delta_{T}^{k}=\begin{bmatrix}\delta_{N}^{j}\cr\delta_{N}^{\ell}\end{bmatrix}$; (17)
  • ii)
    the row vector $(\sum_{\ell=1}^{d}\delta^{i_{\ell}}_{N})^{\top}\left(\sum_{\tau=0}^{N-1}(L_{tot}^{d})^{\tau}\right)$ has (all) positive entries.
Proof.
Problem 1 is solvable if and only if conditions i) and ii) in Proposition 10 are verified for all the BCNs compatible with the data.
In order to check condition i) of Proposition 10, we need to identify from data if each state in ${\mathcal{X}}_{s}$ has (at least) one successor in ${\mathcal{X}}_{s}$. This amounts to verifying if for every $\delta^{j}_{N}\in{\mathcal{X}}_{s}$ there exists $k\in[1,T]$ and $\delta^{\ell}_{N}\in{\mathcal{X}}_{s}$ such that
$\begin{bmatrix}X_{p}\cr X_{f}\end{bmatrix}\delta_{T}^{k}=\begin{bmatrix}\delta_{N}^{j}\cr\delta_{N}^{\ell}\end{bmatrix}$ (18)
which is exactly i) in the theorem statement. If this is the case, an input that allows the transition is
${\bf u}_{j}\triangleq U_{p}\delta_{T}^{k}.$ (19)
On the other hand, by Proposition 6, condition ii) in Proposition 10 is trivially equivalent to condition ii) in the theorem. ∎
Theorem 12 establishes data-based conditions for the solvability of Problem 1; however, it does not yield an explicit solution, which is instead provided in Algorithm 2, below. In Algorithm 2 we first check condition i) of Theorem 12. If such condition is not verified, the algorithm stops and provides a negative outcome. If, on the other hand, the first check is successful, it performs the reachability check required by condition ii) from each state in ${\mathcal{X}}_{u}$ according to the same logic as in Algorithm 1.
Algorithm 2 returns, in Step 1, an input that keeps states already in $\mathcal{X}_{s}$ within $\mathcal{X}_{s}$, and in Step 2, an input that drives states in $\mathcal{X}_{u}$ closer to $\mathcal{X}_{s}$. Such inputs are used to build a state feedback matrix $K$ that solves the problem.
Algorithm 2 Safe control
Input: - The data matrices $X_{p},X_{f}$ and $U_{p}$;
     - The set ${\mathcal{X}}_{s}=\{\delta^{i_{1}}_{N},\delta^{i_{2}}_{N},\dots,\delta^{i_{d}}_{N}\}$.
Output: - Is Problem 1 solvable? Yes/No.
     - A state feedback matrix $K$ that solves the
      problem.
Initialization: $({\bf u}_{1}^{1},{\bf u}_{2}^{1},\dots,{\bf u}_{N}^{1})=(\mathbbold{0},\mathbbold{0},\dots,\mathbbold{0})$,
$({\bf u}_{1}^{2},{\bf u}_{2}^{2},\dots,{\bf u}_{N}^{2})=(\mathbbold{0},\mathbbold{0},\dots,\mathbbold{0})$.
  1. 1.
    • Set $\mathcal{S}_{0}=\emptyset$.
    • for $k\in[1,T]$, do
        if
      $\begin{bmatrix}X_{p}\cr X_{f}\end{bmatrix}\delta_{T}^{k}=\begin{bmatrix}\delta_{N}^{r}\cr\delta_{N}^{q}\end{bmatrix}$,
      with $\delta^{r}_{N}\in{\mathcal{X}}_{s}\setminus{\mathcal{S}}_{0}$ and $\delta^{q}_{N}\in{\mathcal{X}}_{s}$, then
      ${\mathcal{S}}_{0}\leftarrow{\mathcal{S}}_{0}\cup\{\delta^{r}_{N}\}$ and ${\bf u}_{r}^{1}\triangleq U_{p}\delta_{T}^{k}$
@@ 6 Data-driven safe control
      end for
      if
      $\mathcal{S}_{0}\not\equiv\mathcal{X}_{s}$, then
          return (‘No’, ‘None’).
      end if
  2. 2.
    • Apply Algorithm 1 with respect to the set $\mathcal{X}_{s}$.
      Let ($({\bf u}_{1}^{2},{\bf u}_{2}^{2},\dots,{\bf u}_{N}^{2})$, $\mathcal{S}^{*}$) be the corresponding outcome.
    • for $i\in[1,N]$, do
           ${\bf u}_{i}={\bf u}_{i}^{1}+{\bf u}_{i}^{2}$
      end for
    • Set $K=\begin{bmatrix}{\bf u}_{1}&\dots&{\bf u}_{N}\end{bmatrix}$.
  3. 3.
    if $\mathcal{S}^{*}=\mathcal{L}_{N}$, then
        return (‘Yes’, $K$)
    else,
      return (‘No’, ‘None’)
    end if
Example 13, below, explores on a small size network the results of this section pertaining data-driven safe control.
@@
$L$ $=$ $\left[\begin{array}[]{c|c|c}L_{1}&L_{2}&L_{3}\end{array}\right]=\big[\delta_{7}^{4}\ \delta_{7}^{2}\ \delta_{7}^{2}\ \delta_{7}^{5}\ \delta_{7}^{2}\ \delta_{7}^{7}\ \delta_{7}^{5}\ |\$
@@
The equilibrium points of this BCN are $\{\delta_{7}^{2},\delta_{7}^{3},\delta_{7}^{6},\delta_{7}^{7}\}$. We perform a single ($r=1$) offline experiment in the time interval $[0,T]$, with $T=12$, and we collect the following data:
@@
$\tilde{L}$ $=$ $\left[\begin{array}[]{c|c|c}\tilde{L}_{1}&\tilde{L}_{2}&\tilde{L}_{3}\end{array}\right]=\big[\delta_{7}^{4}\ \delta_{7}^{2}\ \delta_{7}^{2}\ *\ *\ *\ *\ |\$
@@
where the symbol * stands for an arbitrary vector (in ${\mathcal{L}}_{7}$). Since
$L_{tot}^{d}=\begin{bmatrix}\delta^{4}_{7}+\delta^{6}_{7}+\delta^{7}_{7}&\ \ \delta^{2}_{7}&\ \ \delta^{2}_{7}+\delta^{3}_{7}&\ \ \delta^{3}_{7}&\ \ \delta^{1}_{7}&\ \delta^{5}_{7}&\ \delta^{6}_{7}+\delta^{7}_{7}\end{bmatrix}$
we obtain that the set of the equilibrium points compatible with all the BCNs in $\mathcal{B}_{d}$ is $\mathcal{X}_{e}^{d}=\{\delta_{7}^{2},\delta_{7}^{3},\delta_{7}^{7}\}.$
Assume now that we want to implement a state feedback safe control with respect to the unsafe set $\mathcal{X}_{u}=\{\delta_{7}^{3},\delta_{7}^{4},\delta_{7}^{7}\}$. We need to understand if the data are informative for safe control with respect to $\mathcal{X}_{u}$, or, equivalently, if conditions i) and ii) of Theorem 12 are satisfied. To this end, we can apply Algorithm 2, where in Step 1 we can readily verify that for every state in $\mathcal{X}_{s}=\mathcal{L}_{7}\setminus\mathcal{X}_{u}=\{\delta_{7}^{1},\delta_{7}^{2},\delta_{7}^{5},\delta_{7}^{6}\}$ there exists a transition ending in a state that is still inside $\mathcal{X}_{s}$, namely condition i) of Theorem 12 holds. In Step 2, we instead examine condition ii) by employing Algorithm 1 to determine the reachability of $\mathcal{X}_{s}$ in every BCN in $\mathcal{B}_{d}$. We obtain: $\mathcal{S}_{0}=\mathcal{X}_{s}$, $\mathcal{S}_{1}=\{\delta_{7}^{3},\delta_{7}^{7}\}$, and $\mathcal{S}_{2}=\{\delta_{7}^{4}\}$, and hence $\mathcal{S}^{*}=\mathcal{L}_{7}$. This means that also condition ii) of Theorem 12 is satisfied, and thus the data are informative for safe control with respect to $\mathcal{X}_{u}$. Algorithm 2 also provides the state feedback matrix $K=\begin{bmatrix}\delta_{3}^{2}&\delta_{3}^{1}&\delta_{3}^{1}&\delta_{3}^{3}&\delta_{3}^{3}&\delta_{3}^{2}&\delta_{3}^{3}\end{bmatrix}$.
@@ 7 Data-driven output regulation
We now address the second problem considered in this paper, namely output regulation, which consists in designing a control input such that the output of a BCN becomes constant and equal to a desired value, after a finite number of steps. More specifically, we pursue this goal by designing a control action in the form of a state feedback.
Before addressing the problem in a data-driven framework, we review its definition and model-based solution (see, e.g., Fornasini & Valcher (2013a, 2014)).
@@
Given a BCN (6), the output regulation problem to ${\bf y}^{*}\in{\mathcal{L}}_{P}$ by state feedback is said to be solvable if there exists $K\in\mathcal{L}_{M\times N}$ such that for every ${\bf x}(0)\in\mathcal{L}_{N}$, ${\bf y}(t)={\bf y}^{*},\forall t\geq\tau$, $\exists\tau\in\mathbb{Z}_{+}$, when ${\bf u}(t)=K{\bf x}(t).$
@@
The first step of the model-based solution consists in finding all the states that generate the desired output $\bf y$$^{*}$, namely the states belonging to the set $\mathcal{X}({\bf y}^{*})\triangleq\{\delta_{N}^{j}:H\delta_{N}^{j}={\bf y}^{*}\}$. Once we have determined this set, we have to ensure that, under some state feedback law ${\bf u}(t)=K{\bf x}(t),\exists K\in\mathcal{L}_{M\times N}$, the state evolution of the resulting BN, described by
${\bf x}(t+1)=L_{K}{\bf x}(t),$ (25)
where $L_{K}\triangleq L\ltimes K\ltimes\Phi_{N}\in\mathcal{L}_{N\times N},$ and $\Phi_{N}$ is the power-reducing matrix Cheng et al. (2011), satisfying $\Phi_{N}{\bf x}(t)={\bf x}(t)\ltimes{\bf x}(t)$, is eventually constrained within $\mathcal{X}({\bf y}^{*})$. This imposes that all limit cycles (in particular, equilibrium points) of the BN (25) have states inside $\mathcal{X}({\bf y}^{*})$.
Since the effect of state feedback on a BCN is that of selecting, for each state, only one of its possible successors, corresponding to a specific choice of the input, each limit cycle appearing in the BN (25) corresponds to a periodic state trajectory of the BCN (6). On the other hand, the periodic state trajectories of the BCN (6) coincide with the cycles appearing in the digraph ${\mathcal{D}}(L_{tot})$. So, if we denote by $\mathcal{C}({\bf y}^{*})$ the set of all cycles in ${\mathcal{D}}(L_{tot})$ whose nodes correspond to states belonging to $\mathcal{X}({\bf y}^{*})$, we first need to verify that $\mathcal{C}({\bf y}^{*})\neq\emptyset$, and then
@@
that it is globally reachable, meaning that the union of the basins of attraction of the states in $\mathcal{C}({\bf y}^{*})$ covers the whole state space $\mathcal{L}_{N}$.
@@
In the following proposition, we formalize the necessary and sufficient conditions for solving the output regulation problem by state feedback in a model-based set-up.
@@
The data-driven version of the output regulation problem is given below.
Problem 2.
Given the data $({\bf x}_{d},{\bf u}_{d},{\bf y}_{d})$, and some desired value ${\bf y}^{*}\in{\mathcal{L}}_{P}$ for the output, determine (if possible) a state feedback matrix $K\in\mathcal{L}_{M\times N}$ such that, by applying ${\bf u}(t)=K{\bf x}(t)$, for every initial condition ${\bf x}(0)$, the output trajectories of all the BCNs compatible with the data satisfy ${\bf y}(t)={\bf y}^{*}$ for every $t\geq\tau,\exists\tau\in\mathbb{Z}_{+}$.
When Problem 2 is solvable, we will say that the data $({\bf x}_{d},{\bf u}_{d},{\bf y}_{d})$ are informative for output regulation to ${\bf y}^{*}\in{\mathcal{L}}_{P}$ by state feedback.
In order to solve Problem 2, we first determine from data the set of states that generate the desired output $\bf y$$^{*}$ in all BCNs in ${\mathcal{B}}_{d}$, i.e.,
$\mathcal{X}^{d}({\bf y}^{*})\!=\!\left\{\delta_{N}^{j}\in{\mathcal{L}}_{N}:\exists k\in[1,T]\ \text{s.t.}\begin{bmatrix}X_{p}\cr Y_{p}\end{bmatrix}\delta_{T}^{k}=\!\begin{bmatrix}\delta_{N}^{j}\cr{\bf y}^{*}\end{bmatrix}\right\}\!$.
To identify this set, we can proceed as in Algorithm 3.
Algorithm 3 Construction of $\mathcal{X}^{d}({\bf y}^{*})$
@@
Output: The set $\mathcal{X}^{d}({\bf y}^{*})$ of all states of the BCNs in
      ${\mathcal{B}}_{d}$ that generate the desired output $\bf y$$^{*}$.
Initialization: Set $\mathcal{X}^{d}({\bf y}^{*})=\emptyset$.
  1. 1.